OpenStudy (anonymous):

3/4-1/2n=5/8 solve for n

OpenStudy (anonymous):

3/4 = 6/8 Move 6/8 to the right side of the equation. 5/8 - 6/8 = - 1/2n -1/8 = -1/2n Since both have negative signs, they cancel out. 1/8 = 1/2n Double the right side to solve for n. n = 2/8 = 1/4